Twice the sum of two numbers is 32. Twice their difference is 4. Find the numbers.
Answer by checkley75(3666)

2(X+Y)=32
X-Y=4 OR X=Y+4
NOW SUBSTITUTE (Y+4) FOR X IN THE OTHERE EQUATON & SOLVE FOR Y.
2(Y+4+Y)=32
2(2Y+4)=32
4Y+8=32
4Y=32-8
4Y=24
Y=24/4
Y=6 ANSWER.
X-6=4
X=4=6
X=10 ANSWER.
PROF
2(10+6)=32
32=32